OpenStudy (anonymous):

Solve the equation for t and express your answer in terms of base-10 logarithms. 4^(√2t+5) =11 t=?

OpenStudy (irishboy123):

log both sides first. if they are equal, their logs should be equal too. use log10 if that is what the question requires. trick is to remember that log (a^ b) is b*log(a)

OpenStudy (anonymous):

irish when i put it in how would i put the ^(√2t+5)

OpenStudy (jdoe0001):

\(\Large \bf 4^{\sqrt{2t}+5}=11\quad ?\)

OpenStudy (irishboy123):

that's the b in the hint i gave. \[\log (x ^{y}) = y \log (x)\] in this schematic, it's the 'y'

OpenStudy (isaiah.feynman):